OpenStudy (campayne):

Use substitution to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) {y = x - 19 2x - y = 27

OpenStudy (3mar):

May I help?

OpenStudy (openstudier777):

Can you help me

OpenStudy (3mar):

Won't be late for any help.

OpenStudy (3mar):

Using substitution: means you solve one equation for a certain variable, the sub with it in the second eq. ok?

OpenStudy (campayne):

right

OpenStudy (3mar):

Good! Can you do that for the first one?

OpenStudy (campayne):

well i dont know how

OpenStudy (3mar):

y = x - 19 This function is already solved for y. Can you sub, with it in the 2nd eq.?

OpenStudy (campayne):

whoa how is it already solved for y?

OpenStudy (3mar):

Y is at one side and the other terms in at the other side!

OpenStudy (3mar):

Is not it?

OpenStudy (campayne):

ohhh i see yes you're right. but idk how to get it that way for the second equation

OpenStudy (3mar):

We don't need that for the 2nd one. Just sub. with the 1st into the 2nd as it is. i.e replace every y in the 2nd eq. with (x-19) got it?

OpenStudy (campayne):

not quite

OpenStudy (3mar):

2nd eq: 2x - y = 27 it will be: 2x-(x-19)=27 clear now?

OpenStudy (3mar):

5 minute please Sorry

OpenStudy (campayne):

its okay, and yes i could figure out that much but what do i do once i have that new equation

OpenStudy (3mar):

Sorry for late.

OpenStudy (3mar):

and Thank you for the medal!

OpenStudy (3mar):

@CamPayne If you are ready to proceed, kindly let me know!

OpenStudy (campayne):

oh its okay! i got it already

OpenStudy (3mar):

Sure!?

OpenStudy (3mar):

I am so sorry that I let you wait! It was an emergency call.

OpenStudy (campayne):

all good bro you take that emergency call

OpenStudy (3mar):

So do you need any help now?